I am a four digit number. The number in my hundreds place is two less than the number in my ones. The number in my thousands place is odd. The number in my tens place is twice the number in my thousands and four more than the number in my ones. What number am I ?
Answer by ankor@dixie-net.com(22740)

I am a four digit number.
a=1000 place
b=100
c=10
d=ones
:
Write an equation for each statement:
:
The number in my hundreds place is two less than the number in my ones.
b = d - 2
:
The number in my thousands place is odd.
:
The number in my tens place is twice the number in my thousands
c = 2a
c has to equal 2, or 6; when a = 1 or 3; try a = 3 and c = 6
:
and four more than the number in my ones.
c = d + 4
6 = d + 4
d = 2
then
b = d - 2
b = 2 - 2
b = 0
:
we now have 3 0 6 2, see if that checks out